# Supp_Appendix_v4

### v3

@@ Abstract
In this paper we study a class of weighted estimands, which we define as parameters that can be expressed as weighted averages of the underlying heterogeneous treatment effects. The popular ordinary least squares (OLS), two-stage least squares (2SLS), and two-way fixed effects (TWFE) estimands are all special cases within our framework. Our focus is on answering two questions concerning weighted estimands. First, under what conditions can they be interpreted as the average treatment effect for some (possibly latent) subpopulation? Second, when these conditions are satisfied, what is the upper bound on the size of that subpopulation, either in absolute terms or relative to a target population of interest? We argue that this upper bound provides a valuable diagnostic for empirical research. When a given weighted estimand corresponds to the average treatment effect for a small subset of the population of interest, we say its internal validity is low. Our paper develops practical tools to quantify the internal validity of weighted estimands.
@@ 1   Introduction
The second and primary aim of this paper is to quantify the degree of validity of $\mu(a,\tau_{0})$ as a causal parameter. To do this, we characterize the size, and the size relative to $W_{0}$, of subpopulations $W^{*}$ associated with the estimand in (1.1). More plainly, we ask how large $\mathbb{P}(W^{*}=1)$ and $\mathbb{P}(W^{*}=1\mid W_{0}=1)$ can be in the representation $\mu(a,\tau_{0})=\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]$. If these probabilities can be large, the estimand corresponds to the average treatment effect for a (relatively) large subpopulation, and when they are small, it corresponds to the average effect for a (relatively) small number of units. If $\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]$ is the target parameter, we interpret a large value of $\mathbb{P}(W^{*}=1\mid W_{0}=1)$ as evidence of a high degree of internal validity of $\mu(a,\tau_{0})$ with respect to the target. If $\mathbb{P}(W^{*}=1)$, the corresponding marginal probability, is large, we say that $\mu(a,\tau_{0})$ is highly representative of the underlying population.
The answer to our questions about subpopulation existence and size depends on the information we have about the CATE function, $\tau_{0}$. Specifically, in one case, we may want to know whether $\mu(a,\tau_{0})$ can be written as $\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]$ for any choice of $\tau_{0}$, or without any knowledge of this function. If this is the case, then we know that the interpretation of $\mu(a,\tau_{0})$ as a causal parameter is robust to heterogeneous treatment effects of any form, including the most adversarial CATE functions. We can also answer the second question about the maximum values of $\mathbb{P}(W^{*}=1)$ and $\mathbb{P}(W^{*}=1\mid W_{0}=1)$ without needing to estimate or know the structure of the CATEs. In a second case, we may want to know how representative $\mu(a,\tau_{0})$ is given knowledge of the CATE function. While the resulting maximum values of $\mathbb{P}(W^{*}=1)$ and $\mathbb{P}(W^{*}=1\mid W_{0}=1)$ are less useful as measures of robustness than in the first case—after all, if the researcher knows or estimates the entire CATE function, they can as well report any average of $\tau_{0}(X)$ that may be relevant—we consider this problem to be of independent theoretical interest. Additionally, if the researcher estimates and compares the maximum values of $\mathbb{P}(W^{*}=1)$ or $\mathbb{P}(W^{*}=1\mid W_{0}=1)$ in both cases, they can evaluate the importance of treatment effect heterogeneity for the interpretation of $\mu(a,\tau_{0})$ in a given application.
@@
Besides theoretical interest, we argue that $\mathbb{P}(W^{*}=1)$ and $\mathbb{P}(W^{*}=1\mid W_{0}=1)$ are a practically relevant measure that may be appealing to applied researchers. First, as stated above, our initial results demonstrate that two commonly invoked criteria for weighted estimands—that they lack negative weights and that they lie in the convex hull of CATE values—are necessary and sufficient (under different assumptions) for the existence of their causal representation, that is, for $\mathbb{P}(W^{*}=1)$ and $\mathbb{P}(W^{*}=1\mid W_{0}=1)$ to be strictly positive. This suggests that researchers invoking these criteria are (indirectly) interested in whether their estimands can be represented as an average treatment effect for some subpopulation. If this is the case, it makes sense to better understand this implicit subpopulation, similar to how it is standard practice in instrumental variables settings to study the subpopulation of compliers. Relatedly, even though our main results concern subpopulation size, we also show that the distribution of covariates in the implicit subpopulation is identified. Thus, practitioners can examine whether this subpopulation has similar characteristics as the entire population, and report the associated sample statistics.
Second, we argue that when $\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]$ is the parameter of interest, it is reassuring for $\mathbb{P}(W^{*}=1\mid W_{0}=1)$ to be large. For one, related claims have been made by other researchers. Mogstad and Torgovitsky (2024) argue that “[t]arget parameters that reflect larger subpopulations of the population of interest are more interesting than those that reflect smaller and more specific subpopulations.” In a setting with multiple instrumental variables, van ’t Hoff, Lewbel, and Mellace (2024) suggest that the largest subpopulation of compliers is generally more interesting than other complier subpopulations. However, we also formalize this claim and show how to construct bounds on $\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]$ that only depend on $\mathbb{P}(W^{*}=1\mid W_{0}=1)$, $\mu(a,\tau_{0})$, and a support restriction. The bounds are easy to compute and collapse to a point as $\mathbb{P}(W^{*}=1\mid W_{0}=1)$ approaches 1. Indeed, large values of $\mathbb{P}(W^{*}=1)$ and $\mathbb{P}(W^{*}=1\mid W_{0}=1)$, our primary measures of interest, guarantee that the weighted estimand is not “too different” from $\mathbb{E}[Y(1)-Y(0)]$ and $\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]$.
@@
This paper is related to a large literature studying weighted average representations of common estimands, including ordinary least squares (OLS), 2SLS, and TWFE in additive linear models. Some of the contributions to this literature include Angrist (1998), Humphreys (2009), Aronow and Samii (2016), Słoczyński (2022), Chen (2024), and Goldsmith-Pinkham, Hull, and Kolesár (2024) for OLS; Imbens and Angrist (1994), Angrist and Imbens (1995), Kolesár (2013), Słoczyński (2020), and Blandhol, Bonney, Mogstad, and Torgovitsky (2022) for 2SLS; and de Chaisemartin and D’Haultfœuille (2020), Goodman-Bacon (2021), Sun and Abraham (2021), Athey and Imbens (2022), Caetano and Callaway (2023), Borusyak, Jaravel, and Spiess (2024), and Callaway, Goodman-Bacon, and Sant’Anna (2024) for TWFE.
@@ 2   Motivating Example
Following Angrist (1998), we can establish that $\beta_{\text{OLS}}$, the coefficient on $D$ in the linear projection of $Y$ on $(1,D,X)$, satisfies the representation in (1.1). The following proposition summarizes Angrist’s (1998) result.
@@
Proposition 2.1 implies that we can write $\beta_{\text{OLS}}$ as a weighted estimand satisfying the representation in (1.1) where $a(X)=p(X)(1-p(X))$ and $\tau_{0}(X)=\mathbb{E}[Y(1)-Y(0)\mid X]$. Here we implicitly set $W_{0}=1$ with probability 1. Thus, the regression coefficient $\beta_{\text{OLS}}$ is a weighted average of CATEs whose weights are $p(X)(1-p(X))$. Note that $\beta_{\text{OLS}}=\text{ATE}\coloneqq\mathbb{E}[Y(1)-Y(0)]$ if and only if $a(X)$ and $\tau_{0}(X)$ are uncorrelated, which is the case, for example, when $p(X)$ or $\tau_{0}(X)$ is constant.
@@ 3.1   Preliminaries
[…]
@@ 3.2   Regular Subpopulations
For convenience, we will abbreviate this as “$W^{*}$ is a regular subpopulation of $W_{0}$”. We denote the set of regular subpopulations of $W_{0}$ as
$\text{SP}(W_{0})$ =$=\{W^{*}\in\{0,1\}:W^{*}\text{ is a regular subpopulation of }W_{0}\}.$
@@ 3.3   Existence of a Causal Representation for Weighted Estimands
For example, if $\tau_{0}$ is constant in $X$, then any weighted estimand satisfying (1.1) equals $\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]$, the average treatment effect within population $\{W_{0}=1\}$. This is the case even when the sign of weight function $a(X)$ varies with $X$. However, if $\tau_{0}$ is nonconstant, the existence of causal representations will depend on the weight function $a(X)$. Among other cases, we will consider the case where no restrictions are placed on function $\tau_{0}$, and in this case, the existence of a causal representation of $\mu(a,\tau_{0})$ will require the sign of $a(X)$ to be constant.
@@
$\mathcal{W}(a;W_{0},\mathcal{T})$ =$=\{W^{*}\in\text{SP}(W_{0}):\mu(a,\tau_{0})=\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]\text{ for all }\tau_{0}\in\mathcal{T}\}.$
This is the set of regular subpopulations of $W_{0}$ such that the estimand $\mu(a,\tau_{0})=\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]$ for all $\tau_{0}$ functions in the set $\mathcal{T}$. If the set $\mathcal{W}(a;W_{0},\mathcal{T})$ is empty, then estimand $\mu(a,\tau_{0})$ cannot be written as an average treatment effect over a regular subpopulation of $W_{0}$ uniformly in $\tau_{0}\in\mathcal{T}$. We use this set to formally define a notion of uniform causal representation.
@@ 3.3.3   Existence in Intermediate Cases
$\mathcal{T}_{\text{lin}}$ =$=\{\tau_{0}\in\mathcal{T}_{\text{all}}:\tau_{0}(X)=c+d^{\prime}X:(c,d)\in\mathbb{R}^{1+d_{X}}\}.$
@@
The above proposition shows that placing restrictions on $\mathcal{T}$ may remove the requirement that $a(X)\geq 0$ for the existence of a uniform causal representation for an estimand. In particular, the requirement here is that $\frac{\mathbb{E}[a(X)X\mid W_{0}=1]}{\mathbb{E}[a(X)\mid W_{0}=1]}$ lies in the convex hull of the support of $X$ given $W_{0}=1$. When $X$ is scalar, this consists of an interval. This condition does not require $a(X)$ be nonnegative. For example, if $\text{supp}(X)=\{0,1,2\}$ and $W_{0}=1$ almost surely, then any combination of values of $(a(0),a(1),a(2))$ such that $\mathbb{E}[a(X)X]/\mathbb{E}[a(X)]\in[0,2]=\text{conv}(\text{supp}(X))$ implies a causal representation. Let $\mathbb{P}(X=x)=1/3$ for $x\in\{0,1,2\}$ and $(a(0),a(1),a(2))=(1,-1,1)$. Here units with $X=1$ have a negative weight, but $\mathbb{E}[a(X)X]/\mathbb{E}[a(X)]=1\in[0,2]$, implying that the corresponding weighted estimand has a causal representation uniformly in $\tau_{0}\in\mathcal{T}_{\text{lin}}$. The result is stated for discrete $X$, but an estimand with negative weights can have a causal representation even when $X$ has continuous components.
@@ 4.1   Quantifying Internal Validity Uniformly in $\tau_{0}$
{adjustwidth}
-1.25in-1.25in Refer to caption Note: $X$ is a single continuously distributed covariate with density $f_{X}$.
@@
[…]

### v4

@@ Abstract
In this paper we study a class of weighted estimands, which we define as parameters that can be expressed as weighted averages of the underlying heterogeneous treatment effects. The popular ordinary least squares (OLS), two-stage least squares (2SLS), and two-way fixed effects (TWFE) estimands are all special cases within our framework. Our focus is on answering two questions concerning weighted estimands. First, under what conditions can they be interpreted as the average treatment effect for some (possibly latent) subpopulation? Second, when these conditions are satisfied, what is the upper bound on the size of that subpopulation, either in absolute terms or relative to a target population of interest? We argue that this upper bound provides a valuable diagnostic for empirical research. When a given weighted estimand corresponds to the average treatment effect for a small subset of the population of interest, we say its internal validity is low. Our paper develops practical tools to quantify the internal validity of weighted estimands. We also apply these tools to revisit a prominent study of the effects of unilateral divorce laws on female suicide.
@@ 1  Introduction
The second and primary aim of this paper is to quantify the degree of validity of $\mu(a,\tau_{0})$ as a causal parameter. To do this, we characterize the size, and the size relative to $W_{0}$, of subpopulations $W^{*}$ associated with the estimand in (1.1). More plainly, we ask how large $\mathbb{P}(W^{*}=1)$ and $\mathbb{P}(W^{*}=1\mid W_{0}=1)$ can be in the representation $\mu(a,\tau_{0})=\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]$. If these probabilities can be large, the estimand corresponds to the average treatment effect for a (relatively) large subpopulation, and when they are small, it corresponds to the average effect for a (relatively) small number of units. If $\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]$ is the target parameter, we interpret a large value of $\mathbb{P}(W^{*}=1\mid W_{0}=1)$ as evidence of a high degree of internal validity of $\mu(a,\tau_{0})$ with respect to the target. [^2] We use the term “internal validity” because it is often associated with the question of whether the probability limit of an estimator is equal to the parameter of interest. If the researcher reports estimates of $\mu(a,\tau_{0})$ when $\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]$ is their ultimate target parameter, then the estimator is internally valid for the target when $\mathbb{P}(W^{*}=1\mid W_{0}=1)=1$. If $\mathbb{P}(W^{*}=1\mid W_{0}=1)<1$, then $\mu(a,\tau_{0})$ will be more informative about the target as the value of $\mathbb{P}(W^{*}=1\mid W_{0}=1)$ increases. If $\mathbb{P}(W^{*}=1)$, the corresponding marginal probability, is large, we say that $\mu(a,\tau_{0})$ is highly representative of the underlying population.
The answer to our questions about subpopulation existence and size depends on the information we have about the CATE function, $\tau_{0}$. Specifically, in one case, we may want to know whether $\mu(a,\tau_{0})$ can be written as $\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]$ for any choice of $\tau_{0}$, or without any restrictions on this function. If this is the case, then we know that the interpretation of $\mu(a,\tau_{0})$ as a causal parameter is robust to heterogeneous treatment effects of any form, including the most adversarial CATE functions. We can also answer the second question about the maximum values of $\mathbb{P}(W^{*}=1)$ and $\mathbb{P}(W^{*}=1\mid W_{0}=1)$ without needing to estimate or know the structure of the CATEs. In a second case, we may want to know how representative $\mu(a,\tau_{0})$ is given knowledge of the CATE function. While the resulting maximum values of $\mathbb{P}(W^{*}=1)$ and $\mathbb{P}(W^{*}=1\mid W_{0}=1)$ are less useful as measures of robustness than in the first case—after all, if the researcher knows or estimates the entire CATE function, they can as well report any average of $\tau_{0}(X)$ that may be relevant—we consider this problem to be of independent theoretical interest. Additionally, if the researcher estimates and compares the maximum values of $\mathbb{P}(W^{*}=1)$ or $\mathbb{P}(W^{*}=1\mid W_{0}=1)$ in both cases, they can evaluate the importance of treatment effect heterogeneity for the interpretation of $\mu(a,\tau_{0})$ in a given application.
@@
Implications for Empirical Practice
Besides theoretical interest, we argue that $\mathbb{P}(W^{*}=1)$ and $\mathbb{P}(W^{*}=1\mid W_{0}=1)$ represent a valuable diagnostic for empirical research. First, as stated above, our initial results demonstrate that two popular criteria for weighted estimands—that they lack negative weights and that they lie in the convex hull of CATE values—are necessary and sufficient (under different assumptions) for the existence of their causal representation, that is, for $\mathbb{P}(W^{*}=1)$ and $\mathbb{P}(W^{*}=1\mid W_{0}=1)$ to be strictly positive. This suggests that researchers invoking these criteria are (indirectly) interested in whether their estimands can be represented as an average treatment effect for some subpopulation. If this is the case, it makes sense to better understand this implicit subpopulation, similar to how it is standard practice in instrumental variables settings to study the subpopulation of compliers. Relatedly, even though our main results concern subpopulation size, we also show that the distribution of covariates in the implicit subpopulation is identified. Thus, practitioners can examine whether this subpopulation has similar characteristics as the entire population, and report the associated sample statistics.
Second, we argue that when $\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]$ is the parameter of interest, it is reassuring for $\mathbb{P}(W^{*}=1\mid W_{0}=1)$ to be large. For one, related claims have been made by other researchers. Mogstad and Torgovitsky (2024) argue that “[t]arget parameters that reflect larger subpopulations of the population of interest are more interesting than those that reflect smaller and more specific subpopulations.” In a setting with multiple instrumental variables, van ’t Hoff, Lewbel, and Mellace (2024) suggest that the largest subpopulation of compliers is generally more interesting than other complier subpopulations. However, we also formalize this claim and show how to construct bounds on $\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]$ that only depend on $\mathbb{P}(W^{*}=1\mid W_{0}=1)$, $\mu(a,\tau_{0})$, and a support restriction. The bounds are easy to compute and collapse to a point as $\mathbb{P}(W^{*}=1\mid W_{0}=1)$ approaches 1. Indeed, large values of $\mathbb{P}(W^{*}=1)$ and $\mathbb{P}(W^{*}=1\mid W_{0}=1)$, our primary measures of interest, guarantee that the weighted estimand is not “too different” from $\mathbb{E}[Y(1)-Y(0)]$ and $\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]$. This makes our measures a practical diagnostic tool to evaluate the robustness of weighted estimands to heterogeneous treatment effects of any form.
@@
This paper is related to a large literature studying weighted average representations of common estimands, including ordinary least squares (OLS), 2SLS, and TWFE in additive linear models. Some of the contributions to this literature include Angrist (1998), Aronow and Samii (2016), Słoczyński (2022), Chen (2024), Goldsmith-Pinkham, Hull, and Kolesár (2024), and Humphreys (2025) for OLS; Imbens and Angrist (1994), Angrist and Imbens (1995), Kolesár (2013), Słoczyński (2020), and Blandhol, Bonney, Mogstad, and Torgovitsky (2022) for 2SLS; and de Chaisemartin and D’Haultfœuille (2020), Goodman-Bacon (2021), Sun and Abraham (2021), Athey and Imbens (2022), Caetano and Callaway (2023), Borusyak, Jaravel, and Spiess (2024), and Callaway, Goodman-Bacon, and Sant’Anna (2024) for TWFE.
@@ 2  Motivating Example
Following Angrist (1998), we can establish that $\beta_{\text{OLS}}$, the coefficient on $D$ in the linear projection of $Y$ on $(1,D,X)$, satisfies the representation in (1.1) under a restriction on the propensity score. The following proposition summarizes Angrist’s (1998) result.
@@
Proposition 2.1 implies that we can write $\beta_{\text{OLS}}$ as a weighted estimand satisfying the representation in (1.1) where a(X)=p(X)(1p(X))a(X)=p(X)(1-p(X)) and $\tau_{0}(X)=\mathbb{E}[Y(1)-Y(0)\mid X]$. Here we implicitly set $W_{0}=1$ with probability 1. Thus, the regression coefficient $\beta_{\text{OLS}}$ is a weighted average of CATEs whose weights are p(X)(1p(X))p(X)(1-p(X)). Note that $\beta_{\text{OLS}}=\text{ATE}\coloneqq\mathbb{E}[Y(1)-Y(0)]$ if and only if $a(X)$ and $\tau_{0}(X)$ are uncorrelated, which is the case, for example, when $p(X)$ or $\tau_{0}(X)$ is constant.
@@ 3.1  Preliminaries
Note that $\tau_{0}$ is defined for all values of $X$ such that $w_{0}(X)=\mathbb{P}(W_{0}=1\mid X)>0$. [^3] While $\tau_{0}(X)$ is only defined when $w_{0}(X)>0$, we set $\tau_{0}(X)w_{0}(X)=0$ when $w_{0}(X)=0$. Throughout this paper, we assume that $\mathbb{P}(W_{0}=1)>0$, so that this subpopulation has a positive mass, which avoids technical issues associated with conditioning on zero-probability events.
@@ 3.2  Regular Subpopulations
For convenience, we will abbreviate this as $W^{*}$ is a regular subpopulation of $W_{0}$”. We denote the set of regular subpopulations of $W_{0}$ as
[…]
@@ 3.3  Existence of a Causal Representation for Weighted Estimands
For example, if $\tau_{0}$ is constant in $X$, then any weighted estimand satisfying (1.1) equals $\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]$, the average treatment effect within population $\{W_{0}=1\}$. This is the case even when the sign of weight function $a(X)$ varies with $X$. However, if $\tau_{0}$ is nonconstant, the existence of causal representations will depend on the weight function $a(X)$. Among other cases, we will consider the case where no restrictions are placed on function $\tau_{0}$. In this case, the existence of a causal representation of $\mu(a,\tau_{0})$ will require the sign of $a(X)$ to be constant.
@@
$\mathcal{W}(a;W_{0},\mathcal{T})$ $=\{W^{*}\in\text{SP}(W_{0}):\mu(a,\tau_{0})=\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]\text{ for all }\tau_{0}\in\mathcal{T}\}.$
This is the set of regular subpopulations of $W_{0}$ such that the estimand $\mu(a,\tau_{0})=\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]$ for all functions $\tau_{0}$ in the set $\mathcal{T}$. If the set $\mathcal{W}(a;W_{0},\mathcal{T})$ is empty, then the estimand $\mu(a,\tau_{0})$ cannot be written as an average treatment effect over a regular subpopulation of $W_{0}$ uniformly in $\tau_{0}\in\mathcal{T}$. We use this set to formally define a notion of uniform causal representation.
@@ 3.3.3  Existence in Intermediate Cases
$\mathcal{T}_{\text{lin}}$ $=\{\tau_{0}\in\mathcal{T}_{\text{all}}:\tau_{0}(X)=c+d^{\prime}X:(c,d)\in\mathbb{R}^{1+d_{X}}\}.$
@@
The above proposition shows that placing restrictions on $\mathcal{T}$ may remove the requirement that $a(X)\geq 0$ for the existence of a uniform causal representation for an estimand. In particular, the requirement here is that $\frac{\mathbb{E}[a(X)X\mid W_{0}=1]}{\mathbb{E}[a(X)\mid W_{0}=1]}$ lies in the convex hull of the support of $X$ given $W_{0}=1$. When $X$ is scalar, this consists of an interval. This condition does not require $a(X)$ be nonnegative. For example, if $\text{supp}(X)=\{0,1,2\}$ and $W_{0}=1$ almost surely, then any combination of values of $(a(0),a(1),a(2))$ such that $\mathbb{E}[a(X)X]/\mathbb{E}[a(X)]\in[0,2]=\text{conv}(\text{supp}(X))$ implies a causal representation. Let $\mathbb{P}(X=x)=1/3$ for $x\in\{0,1,2\}$ and (a(0),a(1),a(2))=(1,1,1)(a(0),a(1),a(2))=(1,-1,1). Here units with $X=1$ have a negative weight, but $\mathbb{E}[a(X)X]/\mathbb{E}[a(X)]=1\in[0,2]$, implying that the corresponding weighted estimand has a causal representation uniformly in $\tau_{0}\in\mathcal{T}_{\text{lin}}$. The result is stated for discrete $X$, but an estimand with negative weights can have a causal representation even when $X$ has continuous components.
@@ 4.1  Quantifying Internal Validity Uniformly in $\tau_{0}$
The construction of this subpopulation is illustrated in Figure 1 for the case where $x$ is continuous and where we omit the conditioning on $W_{0}=1$ for simplicity. We seek to maximize $\mathbb{P}(W^{*}=1)=\int\underline{w}^{*}(x)f_{X}(x)\,dx$ with the requirement that $\underline{w}^{*}(x)\leq 1$ (or, equivalently, $\underline{w}^{*}(x)f_{X}(x)\leq f_{X}(x)$) and that $\underline{w}^{*}(x)$ is a multiple of $a(x)$. In the figure, we see that $a_{\max}>1$ and thus the largest multiple of $a(x)$ that is weakly smaller than 1 is illustrated by the gray curve. The area under this curve is precisely $\mathbb{P}(W^{*}=1)$. Note that the area under $f_{X}(x)$ is one, so closer alignment of the gray curve and the density $f_{X}(x)$ corresponds to more representative estimands.
Note: $X$ is a single continuously distributed covariate with density $f_{X}$.
@@
Remark 4.4 (Meaning of internal validity and representativeness).
Suppose we consider $\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]$ to be the target parameter, where $W^{*}=\mathbbm{1}\left(U\leq a(X)/a_{\max}\right)\cdot W_{0}$ is the subpopulation for which $\mu(a,\tau_{0})=\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]$ uniformly in $\tau_{0}$. For example, in the case of the OLS estimand in Section 2, this consists of a subpopulation where the probability of inclusion is proportional to the conditional variance of treatment. If this subpopulation is the target, it would be reasonable to infer that the measure of internal validity of the estimand $\mu(a,\tau_{0})$ is the maximum value of 1. Indeed, this is the case because the estimand can be written as
$\mu(a,\tau_{0})=\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]$
and $\{W^{*}=1\}$ is trivially the largest regular subpopulation of $\{W^{*}=1\}$. This example illustrates how the internal validity of $\mu(a,\tau_{0})$ is entirely dependent on the target estimand. This is appropriate because we associate the term “internal validity” with the question of whether the probability limit of an estimator is equal to the parameter of interest. Thus, the degree of internal validity should naturally depend on the parameter choice. On the other hand, the representativeness of the weighted estimand, as measured by the largest value of $\mathbb{P}(W^{*}=1)$, is independent of the target parameter. In fact, it will be less than one unless $W^{*}=1$ almost surely, or that the weighted estimand actually equals the ATE. If, in addition, we wish to operationalize the concept of external validity in our framework, we need to consider whether the weighted estimand can be written as the average treatment effect within a subpopulation of another, possibly arbitrary population. This extension is left for future work.